OpenStudy (anonymous):

What is the solution to the equation 3 1 + 2x = 2 ?

OpenStudy (anonymous):

Is the equation 3×1+2 x = 2? or maybe 31+2 x = 2?

OpenStudy (anonymous):

*3^1+2x=2

OpenStudy (anonymous):

3^1+2x = 2 2x+3 = 2 2 x = -1 x = -1/2